# Modular transformations of on-shell actions of (root-)$\overline{\text{T}}$ deformed holographic CFTs

### v2

@@ 1 Introduction
In the past two decades, a very interesting irrelevant deformation of two-dimensional conformal field theories (CFTs), known as $\overline{\text{T}}$ deformation, was proposed [1, 2, 3]. One of the most intriguing properties of $\overline{\text{T}}$ deformed CFTs is that one can exactly calculate some quantities including energy spectrum [1, 2, 3] and S-Matrix [2, 4]. [^4] Moreover, the closed form of the Lagrangian of the $\overline{\text{T}}$ deformed theory can be obtained for some models, e.g. [5, 6]. In recent years, various features of the $\overline{\text{T}}$ deformed CFTs were investigated such as Lax connections[7, 8], S-Matrix [4, 9], correlation functions [11, 19, 10, 18, 12, 13, 14, 16, 20, 17, 26, 25, 15, 21, 22, 23, 24, 27], partition function[28, 29, 30, 32, 35, 33, 34, 31], entanglement entropy [36, 37, 38, 39, 41, 42, 40, 43, 55, 49, 50, 51, 52, 48, 46, 54, 53, 44, 45, 47, 56], pseudo entanglement entropy [57, 58], entanglement negativity [59], mutual information [61, 60, 52], entanglement wedge cross section [61, 60, 62, 56] and Action-Complexity [63, 65, 64]. Furthermore, theses deformations were extended to quantum field theories in zero dimension [66], one dimension [67, 68, 69], higher dimensions [71, 5, 70, 72, 73], to non-relativistic quantum field theories [74, 65, 6, 60] as well as to lattice theories [75]. See also [76, 77] which are very nice and comprehensive reviews on the recent developments.
@@
In the context of the AdS/CFT correspondence [86] , there are two different proposals for the holographic dual of $\overline{\text{T}}$ deformed CFTs: [^5] The holographic dual of root-$\overline{\text{T}}$ deformed CFTs was studied in [85]. cut-off proposal [87, 10] (or glue-on proposal [54, 88]) and mixed-boundary-condition proposal [89]. The cut-off proposal is easier to impliment, but it has several limitations[89]. On the other hand, the mixed-boundary-condition proposal is applicable to all holographic theories, regardless of the presence of matters, and the holographic dictionary can be summarized as follows: [77] 666the notation will become clear in a moment.
@@
where $\mu$ is the deformation parameter. The key observation of us is that in the dictionary (1) we should add a proper boundary term to the gravity action [^7] Perhaps this is already hidden written in[89] and the presence of this them is stressed by one of the authors in [55].. The presence of this boundary term is natural considering that the double trace deformation not only shifts the source but also shifts the generating function [90, 91]. In this paper, we compute the on-shell actions of the $\overline{\text{T}}$ deformed theory defined on a torus. The aim is to derive and understand the modular properties of the $\overline{\text{T}}$ deformed torus partition function from the holographic perspective.
A two-dimensional CFT defined on a torus (and other Riemann surfaces) is subject to an important set of consistent conditions known as modular invariance. This requires that the partition function or correlation functions should remain the same, regardless of how the torus is represented. Modular invariance is a powerful constraint, leading to significant consequences such as the Hawking-Page transition [92] and the Cardy formula [93]. It has been discovered in [29] that the deformation parameter should also transform under the modular transformation which means that this kind of transformation can relate the partition functions of two different quantum field theories along the $\overline{\text{T}}$ flow. We find that this property can be easily understood by observing that the deformation parameter $\mu$ has the same dimension of the radial coordinate and undergoes rescaling under a Weyl transformation. Meanwhile, we derive the thermodynamic first law and Smarr relations for the deformed BTZ black hole. Furthermore, we show that the modular property remains preserved even in the CFT with a gravitational anomaly[94], where the holomorphic sector and anti-holomorphic sector have different central charges[95], which is dual to a topologically massive gravity (TMG) theory [96, 97]. As a contrast, we also consider a marginal deformation, the root-$\overline{\text{T}}$ deformation [82, 83] in the mixed-boundary-condition proposal [85], for which the deformation parameter is dimensionless and we find indeed it does not transform under modular transformation.
@@ 2 Mixed-boundary-condition proposal and the deformed on-shell action
The specialty of this double trace deformation is that the operator $\overline{\text{T}}$ depends both on the boundary metric and the energy-stress tensor. Consequently, both of them flow under the deformation according to [89]:
@@
where $\hat{T}_{\alpha\beta}=T_{\alpha\beta}-\gamma_{\alpha\beta}T$, see Appendix A for details. The initial data $(\gamma_{\alpha\beta}^{[0]},\hat{T}^{[0]}_{\alpha\beta})$ defines the seed CFT which generally should not be viewed as the undeformed CFT because from the field theory point of view the $\overline{\text{T}}$ deformation only deforms the theory but not the spacetime background [^8] Do not be confused with the dynamical coordinate transformation interpretation [98, 99] of the $\overline{\text{T}}$ deformation.. The seed CFT is dual to an asymptotic AdS3 solution which can be written in the Fefferman-Graham gauge [100]
@@
We will interchangeably use $\rho_{c}$ and $\mu$ as the deformation parameters.
As pointed out in [55], the proper on-shell action of the $\overline{\text{T}}$ deformed theory should be
@@
Using (2.4) and (2.5) one can show that $\sqrt{\gamma}\text{T$\overline{\text{T}}$ }$ is invariant under the flow so the boundary term in (2.9) can also be written as
@@
$ds^{2}=\frac{dz^{2}}{z^{2}}+\frac{1}{z^{2}}(dv+\bar{L}_{0}(\bar{v})z^{2}d\bar{v})(d\bar{v}+L_{0}(v)z^{2}dv).$ (2.11)
In this case, one can show that the deformed metric $\gamma_{\alpha\beta}^{[\mu]}$ is also flat thus (2.8) is related to $g_{\alpha\beta}^{(0)}$ via the coordinate transformations
@@
where $L_{0}$ and $\bar{L}_{0}$ should be understood as functions of $V$ and $\bar{V}$. The deformed stress tensor is given by (2.5):
@@
where $c=3/(2G_{N})$ [101]. Since we are interested in the modular property of the torus partition function, below we will assume that the boundary manifold is a flat torus with two generic complex periods $\bm{\omega}=\omega_{1}+\text{i}\omega_{2}$ and $\bm{\beta}=\beta_{1}+\text{i}\beta_{2}$ so the possible bulk solutions are global AdS, BTZ black hole and other $SL(2,R)$ black holes. Because AdS3 gravity does not have local degrees of freedom, these solutions are all diffeomorphic equivalent but distinguished by the periodicity. A torus is specified by the modular parameter defined by the ratio of the two complex periods: $\bm{\tau}=\frac{\bm{\beta}}{\bm{\omega}}$. It is well known that the undeformed on-shell actions of these solutions are related to each other via the $SL(2,\mathbb{Z})$ modular transformation. This modular property can be traced back to the diffeomorphic equivalence of bulk solutions. However, it may be broken due to the non-commutativity of the $\overline{\text{T}}$ deformation with the diffeomorphism. Nevertheless, [29] shows that the $\overline{\text{T}}$ deformed torus partition function exhibits a modified modular invariance:
@@
We want to emphasize that during the $\overline{\text{T}}$ deformation process, three distinct CFTs are involved: the seed CFT defined on the seed torus with metric $ds^{2}=dvd\bar{v}$, the deformed CFT defined on the deformed torus with metric $ds^{2}=dVd\bar{V}$, and the undeformed CFT defined on the undeformed torus with metric $d\tilde{s}^{2}=d\tilde{v}d\bar{\tilde{v}}$. The explicit relationships between these three CFTs will be discussed in each of the subsequent examples.
@@ 3 General global AdS
whose metric is then
@@
${L}_{0}=-\frac{\pi^{2}}{\bm{\omega}^{2}},\quad\bar{{L}}_{0}=-\frac{\pi^{2}}{\bar{\bm{\omega}}^{2}}.$ (3.4)
@@ 3.1 $\overline{\text{T}}$ deformation
According to the dictionary (2.13), the $\overline{\text{T}}$ deformed theory is also defined on a torus with a flat metric $ds^{2}=dVd\bar{V}$, which is related to seed metric via
@@
$\bm{\omega}_{\mu}=\bm{\omega}+\bar{L}_{0}\rho_{c}{\bm{\bar{\omega}}}=\bm{\omega}(1-\frac{\pi^{2}\rho_{c}}{\bm{\omega}^{2}}),\quad\bm{\beta}_{\mu}=\bm{\beta}-\frac{\pi^{2}\rho_{c}}{\bm{\omega}^{2}}\bm{\bar{\beta}},$ (3.9)
@@
which is consistent with the field theory calculation of the deformed energy [1, 2, 3]. It may be confusing that the undeformed energy $M_{0}$ depends on $\bm{\omega}_{\mu}$ which depends on the deformation parameter $\rho_{c}$ implicitly via (3.9). However, as we emphasized in last section that when we compare the deformed CFT and the undeformed one we should fix the size of the system. That is why the undeformed energy $M_{0}$ can be obtained directly by setting $\rho_{c}=0$ in (3.11).
@@ 4 The BTZ black hole
An Euclidean rotating BTZ black hole has the standard metric
@@
where its two (complex) periods
$(\phi,t)\sim(\phi+2\pi,t),\quad(\phi,t)\sim(\phi+\theta,t+\beta),$ (4.2)
are related to the inner and outer horizon radii via
@@
For more details refer to Appendix B. Because the effect of the rescaling is like a boost (or more precisely a rotation in the Euclidean signature), the energy and the momentum get transformed accordingly and take a slightly complicated form
@@
The details of the these deviations can be found in Appendix B. It is straightforward to check that the Smarr relation
@@
$S_{BH}=\sqrt{\frac{c\pi\bm{\omega}}{6}{(M+\text{i}J)}{}}+\sqrt{\frac{c\pi\bm{\bar{\omega}}}{6}{(M-\text{i}J)}{}}$ (4.9)
@@ 4.1 The $\overline{\text{T}}$ deformation
The $\overline{\text{T}}$ deformed theory is also defined on a torus with metric $ds^{2}=dVd\bar{V}$ where the deformed coordinates $(V,\bar{V})$ are related to seed metric via (3.7) and (3.8), however now $L_{0}$ and $\bar{L}_{0}$ are given by (4.4). It implies that the two complex periods of the deformed torus are
@@
and the $\overline{\text{T}}$ deformed bulk geometry is described by the corresponding metric (2). The relation (4.10) implies that $\bm{\omega}_{\mu}$ is complex in general and the geometry describes a general deformed $SL(2,\mathbb{Z})$ black hole which we will consider later on. For BTZ solutions, we choose $\bm{\omega}_{\mu}$ to be real. The deformed energy and momentum can be found directly by using (2.15).
The energy and the momentum of the deformed theory are given by
@@
In the limit $\rho_{c}=0$, they reduce to
@@
However, $M_{0}$ and $J_{0}$ are not the undeformed energy and momentum. To see this, for example by setting $J_{0}=J_{\mu}=0$ (i.e. $\bm{\beta}_{\mu}=-\bm{\bar{\beta}}_{\mu}$) we find
@@
When $\bm{\beta}_{\mu}=\text{i}\beta_{\mu}$ is purely imaginary, so are $\bm{\beta}=\text{i}\beta$ and $\bm{\tilde{\beta}}=\text{i}\tilde{\beta}$ then the relation (4.17) is simplified to
@@
This coordinate transformation (4.16) is uniquely fixed by requiring that it does not change the horizon area (black hole entropy) and the momentum which are supposed to be invariant under $\overline{\text{T}}$ flow. Clearly, the undeformed CFT and the deformed CFT have the same spatial dimension. Note that the undeformed CFT and the seed CFT is related by a Weyl transformation:
@@
$=$ $-\frac{cA_{\mu}}{6\pi}\frac{\pi^{2}}{|\bm{\beta}_{\mu}|^{2}}{+}\rho_{c}\frac{cA_{\mu}}{6\pi}\frac{\pi^{4}}{|\bm{\beta}_{\mu}|^{2}}+\mathcal{O}(\rho_{c}^{2}),$

### v3

@@ 1 Introduction
In the past two decades, a very interesting irrelevant deformation of two-dimensional conformal field theories (CFTs), known as $\overline{\text{T}}$ deformation, was proposed [1, 2, 3]. One of the most intriguing properties of $\overline{\text{T}}$ deformed CFTs is that one can exactly calculate some quantities including energy spectrum [1, 2, 3] and S-Matrix [2, 4]. [^4] Moreover, the closed form of the Lagrangian of the $\overline{\text{T}}$ deformed theory can be obtained for some models, e.g. [5, 6]. In recent years, various features of the $\overline{\text{T}}$ deformed CFTs were investigated such as Lax connections[7, 8], S-Matrix [4, 9], correlation functions [11, 19, 10, 18, 12, 13, 14, 16, 20, 17, 26, 25, 15, 21, 22, 23, 24, 27], partition function[28, 29, 30, 31, 33, 36, 34, 35, 32, 37], entanglement entropy [38, 39, 40, 41, 43, 44, 42, 45, 57, 51, 52, 53, 54, 50, 48, 56, 55, 46, 47, 49, 58], pseudo entanglement entropy [59, 60], entanglement negativity [61], mutual information [63, 62, 54], entanglement wedge cross section [63, 62, 64, 58] and Action-Complexity [65, 67, 66]. Furthermore, theses deformations were extended to quantum field theories in zero dimension [68], one dimension [69, 70, 71], higher dimensions [73, 5, 72, 74, 75, 76], to non-relativistic quantum field theories [77, 67, 6, 62] as well as to lattice theories [78]. See also [79, 80] which are very nice and comprehensive reviews on the recent developments.
@@
In the context of the AdS/CFT correspondence [89] , there are two different proposals for the holographic dual of $\overline{\text{T}}$ deformed CFTs: [^5] The holographic dual of root-$\overline{\text{T}}$ deformed CFTs was studied in [88]. cut-off proposal [90, 10] (or glue-on proposal [56, 91]) and mixed-boundary-condition proposal [92]. The cut-off proposal is easier to implement, but it has several limitations[92]. On the other hand, the mixed-boundary-condition proposal is applicable to all holographic theories, regardless of the presence of matters, and the holographic dictionary can be summarized as follows: [80]  [^6] The notation will become clear in a moment.
@@
where $\mu$ is the deformation parameter. The key observation of us is that in the dictionary (1) we should add a proper boundary term to the gravity action (See eq. (2.10))  [^7] Perhaps this is already hidden written in [92] and the presence of this term is stressed by one of the authors in [57].. The presence of this boundary term is natural considering that the double trace deformation not only shifts the source but also shifts the generating function [93, 94]. In this paper, we compute the on-shell actions of the $\overline{\text{T}}$ deformed theory defined on a torus. The aim is to derive and understand the modular properties of the $\overline{\text{T}}$ deformed torus partition function from the holographic perspective. Interestingly, the new boundary term proves to be crucial for the on-shell action to exhibit the appropriate modular property and other desirable characteristics. It is also worth noting that in the cut-off proposal of the $\overline{\text{T}}$ deformation, the presence of a similar term has been suggested [95]. However, in the cut-off proposal, the boundary term is introduced at the cut-off surface instead of the AdS boundary.
A two-dimensional CFT defined on a torus (and other Riemann surfaces) is subject to an important set of consistent conditions known as modular invariance. This requires that the partition function or correlation functions should remain the same, regardless of how the torus is represented. Modular invariance is a powerful constraint, leading to significant consequences such as the Hawking-Page transition [96] and the Cardy formula [97]. It has been discovered in [29] and also in [30] that the deformation parameter should also transform under the modular transformation which means that this kind of transformation can relate the partition functions of two different quantum field theories along the $\overline{\text{T}}$ flow. We find that this property can be easily understood by observing that the deformation parameter $\mu$ has the same dimension of the radial coordinate and undergoes rescaling under a Weyl transformation. Meanwhile, we derive the thermodynamic first law and Smarr relations for the deformed BTZ black hole. Furthermore, we show that the modular property remains preserved even in the CFT with a gravitational anomaly[98], where the holomorphic sector and anti-holomorphic sector have different central charges[99], which is dual to a topologically massive gravity (TMG) theory [100, 101]. As a contrast, we also consider a marginal deformation, the root-$\overline{\text{T}}$ deformation [85, 86] in the mixed-boundary-condition proposal [88], for which the deformation parameter is dimensionless and we find indeed it does not transform under modular transformation.
@@ 2 Mixed-boundary-condition proposal and the deformed on-shell action
The significance of this double trace deformation is that the operator $\overline{\text{T}}$ depends both on the boundary metric and the energy-stress tensor. Consequently, both of them flow under the deformation according to [92]:
@@
where $\hat{T}_{\alpha\beta}=T_{\alpha\beta}-\gamma_{\alpha\beta}T$, see Appendix A for more details. The initial data $(\gamma_{\alpha\beta}^{[0]},\hat{T}^{[0]}_{\alpha\beta})$ defines the seed CFT which generally should not be considered as the undeformed CFT because from the field theory point of view the $\overline{\text{T}}$ deformation only deforms the theory but not the spacetime background [^8] Do not be confused with the dynamical coordinate transformation interpretation [102, 95] of the $\overline{\text{T}}$ deformation.. The seed CFT is dual to an asymptotic AdS3 solution and its metric can be written in the Fefferman-Graham gauge as follows [103]
@@
In the following, we will interchangeably use $\rho_{c}$ and $\mu$ as the deformation parameters.
As pointed out in [95, 57], the proper on-shell action of the $\overline{\text{T}}$ deformed theory should be
@@
Using (2.4) and (2.5), one can show that $\sqrt{\gamma}\text{T$\overline{\text{T}}$ }$ is invariant under the flow, so the boundary term in (2.9) can also be written as
@@
$ds^{2}=\frac{dz^{2}}{z^{2}}+\frac{1}{z^{2}}\left(dv+\bar{L}_{0}(\bar{v})z^{2}d\bar{v}\right)\left(d\bar{v}+L_{0}(v)z^{2}dv\right).$ (2.11)
In this case, one can show that the deformed metric $\gamma_{\alpha\beta}^{[\mu]}$ is also flat thus (2.8) is related to $g_{\alpha\beta}^{(0)}$ via the following coordinate transformations
@@
where $L_{0}$ and $\bar{L}_{0}$ should be understood as functions of $V$ and $\bar{V}$. Moreover, the deformed stress tensor is given by (2.5):
@@
where $c=3/(2G_{N})$ [104]. Since we are interested in the modular property of the torus partition function, below we will assume that the boundary manifold is a flat torus with two generic complex periods $\bm{\omega}=\omega_{1}+\text{i}\omega_{2}$ and $\bm{\beta}=\beta_{1}+\text{i}\beta_{2}$. Therefore, the possible bulk solutions are global AdS, BTZ black hole and other $SL(2,R)$ black holes. Because AdS3 gravity does not have local degrees of freedom, these solutions are all diffeomorphic equivalent but distinguished by the periodicity. A torus is specified by the modular parameter defined by the ratio of the two complex periods: $\bm{\tau}=\frac{\bm{\beta}}{\bm{\omega}}$. It is well known that the undeformed on-shell actions of these solutions are related to each other via the $SL(2,\mathbb{Z})$ modular transformation. This modular property can be traced back to the diffeomorphic equivalence of bulk solutions. However, it may be broken due to the non-commutativity of the $\overline{\text{T}}$ deformation with the diffeomorphism. Nevertheless, [29] shows that the $\overline{\text{T}}$ deformed torus partition function exhibits a modified modular invariance:
@@
We want to emphasize that during the $\overline{\text{T}}$ deformation process, three distinct CFTs are involved: the seed CFT defined on the seed torus with metric $ds^{2}=dvd\bar{v}$, the deformed CFT defined on the deformed torus with metric $ds^{2}=dVd\bar{V}$, and the undeformed CFT defined on the undeformed torus with metric $d\tilde{s}^{2}=d\tilde{v}d\bar{\tilde{v}}$. The explicit relationships among these three CFTs will be discussed in each of the subsequent examples.
@@ 3 General global AdS
whose metric is given by
@@
${L}_{0}=-\frac{\pi^{2}}{\bm{\omega}^{2}},\quad\bar{{L}}_{0}=-\frac{\pi^{2}}{\bar{\bm{\omega}}^{2}},$ (3.4)
@@ 3.1 $\overline{\text{T}}$ deformation
According to the dictionary (2.13), the $\overline{\text{T}}$ deformed theory is also defined on a torus with a flat metric $ds^{2}=dVd\bar{V}$, which is related to the seed metric via
@@
$\bm{\omega}_{\mu}=\bm{\omega}+\bar{L}_{0}\rho_{c}{\bm{\bar{\omega}}}=\bm{\omega}\left(1-\frac{\pi^{2}\rho_{c}}{\bm{\omega}^{2}}\right),\quad\bm{\beta}_{\mu}=\bm{\beta}-\frac{\pi^{2}\rho_{c}}{\bm{\omega}^{2}}\bm{\bar{\beta}},$ (3.9)
@@
which is consistent with the field theory calculation of the deformed energy [1, 2, 3]. It may be confusing that the undeformed energy $M_{0}$ depends on $\bm{\omega}_{\mu}$ which depends on the deformation parameter $\rho_{c}$ implicitly via (3.9). However, as we emphasized in last section when we compare the deformed CFT and the undeformed one, we should fix the size of the system. That is why the undeformed energy $M_{0}$ can be obtained directly by setting $\rho_{c}=0$ in (3.11).
@@ 4 The BTZ black hole
A Euclidean rotating BTZ black hole has the standard metric
@@
where
$(\phi,t)\sim(\phi+2\pi,t),\quad(\phi,t)\sim(\phi+\theta,t+\beta).$ (4.2)
Moreover, its two (complex) periods $\beta$ and $\theta$ are related to the inner and outer horizon radii via
@@
For more details refer to Appendix B. Because the effect of the rescaling is similar to a boost (or more precisely a rotation in the Euclidean signature), the energy and momentum get transformed accordingly and take a slightly complicated form
@@
The details of these derivations can be found in Appendix B. It is straightforward to check that the Smarr relation
@@
$S_{BH}=\sqrt{\frac{c\pi\bm{\omega}}{6}{(M+\text{i}J)}{}}+\sqrt{\frac{c\pi\bm{\bar{\omega}}}{6}{(M-\text{i}J)}{}},$ (4.9)
@@ 4.1 The $\overline{\text{T}}$ deformation
The $\overline{\text{T}}$ deformed theory is also defined on a torus with metric $ds^{2}=dVd\bar{V}$ where the deformed coordinates $(V,\bar{V})$ are related to the seed metric via (3.7) and (3.8), however now $L_{0}$ and $\bar{L}_{0}$ are given by (4.4). It implies that the two complex periods of the deformed torus are
@@
and the $\overline{\text{T}}$ deformed bulk geometry is described by the corresponding metric (2). The relation (4.10) implies that $\bm{\omega}_{\mu}$ is complex in general and the geometry describes a general deformed $SL(2,\mathbb{Z})$ black hole which we will consider later on. For BTZ solutions, we choose $\bm{\omega}_{\mu}$ to be real. The energy and momentum of the deformed theory can be found directly by using (2.15) as follows
@@
In the limit $\rho_{c}\rightarrow 0$, they reduce to
@@
However, $M_{0}$ and $J_{0}$ are not the undeformed energy and momentum. To see this, for example we set $J_{0}=J_{\mu}=0$ (i.e. $\bm{\beta}_{\mu}=-\bm{\bar{\beta}}_{\mu}$), then we find
@@
When $\bm{\beta}_{\mu}=\text{i}\beta_{\mu}$ is purely imaginary, we have $\bm{\beta}=\text{i}\beta$ and $\bm{\tilde{\beta}}=\text{i}\tilde{\beta}$, then the relation (4.17) is simplified to
@@
This coordinate transformation (4.16) is uniquely fixed by requiring that it does not change the horizon area (black hole entropy) and the momentum which are supposed to be invariant under the $\overline{\text{T}}$ flow. Clearly, the undeformed CFT and the deformed CFT have the same spatial dimension. Note that the undeformed CFT and the seed CFT are related by a Weyl transformation:
@@
$=$ $-\frac{cA_{\mu}}{6\pi}\frac{\pi^{2}}{|\bm{\beta}_{\mu}|^{2}}{+}\rho_{c}\frac{cA_{\mu}}{6\pi}\frac{\pi^{4}}{|\bm{\beta}_{\mu}|^{2}}+\mathcal{O}(\rho_{c}^{2})$